### v1

@@ Abstract
In recent years, machine learning models have seen widespread adoption across a broad range of sectors, including high-stakes domains such as healthcare, finance, and law. This growing reliance has raised increasing concerns regarding model interpretability and accountability, particularly as legal and regulatory frameworks place tighter constraints on the use of so-called black-box models in critical applications.
Although interpretable machine learning has attracted substantial attention, systematic evaluations of inherently interpretable models, especially for tabular data, remain relatively scarce and often focus primarily on aggregated performance outcomes.
To address this gap, we present a large-scale comparative evaluation of sixteen inherently interpretable methods, ranging from classical linear models and decision trees to more recent approaches such as Explainable Boosting Machines (EBMs), Symbolic Regression, and Generalized Optimal Sparse Decision Trees (GOSDT). Our study spans 216 real-world tabular datasets and goes beyond aggregate rankings by stratifying performance according to structural dataset characteristics, including dimensionality, sample size, linearity, and class imbalance. In addition, we assess training time and robustness under controlled distributional shifts.
Our results reveal clear performance hierarchies, particularly for regression tasks, where EBMs consistently achieve strong predictive accuracy. At the same time, we show that performance is highly context-dependent: Symbolic Regression and Interpretable Generalized Additive Neural Networks (IGANNs) perform particularly well in non-linear regimes, while GOSDT models exhibit pronounced sensitivity to class imbalance.
Overall, these findings provide practical guidance for practitioners seeking a balance between interpretability and predictive performance, and contribute to a deeper empirical understanding of interpretable modeling for tabular data.
@@ 1 Introduction
The field of Interpretable Machine Learning (IML) has emerged, emphasizing models whose internal mechanisms are transparent and understandable to human users (molnar2020interpretable). This paradigm stands in contrast to the black box nature of complex architectures, such as deep neural networks, which prioritize predictive accuracy at the expense of explanatory insight (breiman2001statistical; rudin2019we).
@@
It is also necessary to distinguish intrinsic interpretability from post-hoc explainability (often termed XAI). Post-hoc methods attempt to explain a complex model after it has been trained; however, because these are merely approximations, they may not accurately reflect how the model actually made a decision. In contrast, intrinsically interpretable models are transparent by design, guaranteeing that the explanation is an exact representation of the model’s internal logic (rudin_stop_2019). For this reason, such techniques are increasingly adopted in domains such as healthcare, finance, and law, where accountability and the understanding of the model decision-making process are essential (doshi2017towards; rudin_stop_2019). This demand is further amplified by evolving regulatory frameworks, such as the European Union’s AI Act and the General Data Protection Regulation (GDPR), which mandate stringent transparency requirements and the right to explanation for high-risk automated systems. In many tabular-data applications, intrinsically interpretable models, such as linear models, sparse decision trees, or symbolic models, are already regarded as strong candidates due to their clarity and ease of validation by domain experts (zeng2017interpretable; la2021contemporary). Despite this growing practical importance, empirical guidance on the comparative performance of interpretable models remains limited.
@@
Most existing benchmarking studies focus on general-purpose machine learning algorithms rather than interpretable ones (fernandez2014we; wainer2016comparison; olson2017pmlb; scholz_comparison_2021). As a result, practitioners often lack systematic evidence about (i) whether some interpretable models tend to perform better than others across diverse real-world datasets, and (ii) how model performance changes as a function of dataset characteristics such as dimensionality, sample size, linearity, or class imbalance. A few recent works examine specific subclasses of interpretable models, such as Generalized Additive Models (GAMs) (kruschel2025challenging), but do not cover the broader landscape of modern IML approaches nor analyze performance variation under different data conditions. Moreover, comparative evaluations are typically conducted on small collections of hand-selected datasets or on synthetic data, which may limit generalizability.
@@
At the same time, the ecosystem of interpretable methods has expanded substantially over the past decade. Techniques such as Explainable Boosting Machines (EBMs) (nori2019interpretml), Interpretable Generalized Additive Neural Networks (IGANNs) (kraus2023interpretable), Generalized Optimal Sparse Decision Trees (GOSDTs) (mctavish2022fast), and modern Symbolic Regression methods (la2021contemporary) provide new modeling options whose empirical properties have not yet been systematically compared. Consequently, users are faced with an increasingly fragmented methodological landscape without clear, data-driven recommendations for model selection.
@@
  • Using the full Penn Machine Learning Benchmark (PMLB) repository (olson2017pmlb), we evaluate 16 interpretable models across 137 regression and 79 classification datasets. Our analysis includes both traditional and recent IML methods with stable, well-maintained implementations (9 regression models and 7 classification models), assessing both predictive performance and training time. Our experimental design includes 4-fold cross-validation, systematic hyperparameter optimization, and rigorous preprocessing, resulting in over 14,000 model evaluations and more than 700,000 individual training runs.
  • Beyond aggregated results, we investigate how model performance varies with key dataset characteristics, including structural properties (e.g., number of features and samples), degree of linearity, and class imbalance (maciel_measuring_2016; lorena_how_2020). This stratified analysis provides deeper insights into the conditions under which different interpretable models excel or underperform, offering practical guidance for both researchers and practitioners.
@@
Our empirical analysis reveals several important patterns. For regression tasks, clear performance hierarchies emerge under different data regimes: methods such as EBM and Symbolic Regression (SR) consistently excel on non-linear or low-sample problems, whereas regularized models like LASSO and Generalized Linear Models (GLMs) demonstrate superior robustness against distributional shifts and high dimensionality compared to standard Linear Regression (LR). In contrast, for classification, performance differences between models are less sharply stratified: most methods behave similarly across data conditions, making it more difficult to identify a preferable model a priori. This weaker stratification is partly attributable to the smaller number of available classification datasets in PMLB, but it also suggests that classification performance may be inherently less sensitive to the dataset metrics considered here.
The remainder of this paper is organized as follows. Section 2 reviews related work on benchmarking with a focus on interpretable models. Section 3 details the experimental setup, including the selection of machine learning models and dataset preprocessing. Section 4 specifies the evaluation protocol, defining the cross-validation design, generalization scenarios, and performance metrics. Section 5 describes the dataset metrics used for stratification and the statistical procedures for model ranking. Sections 6 and 7 present and analyze the empirical results for the regression and classification tasks. Section 8 relates our results to existing work and discusses their implications. Section 9 concludes the paper and highlights directions for future research.
@@ 2 Related Work
A number of studies in the literature have conducted comparative analyses of machine learning models (fernandez2014we; wainer2016comparison; olson2017pmlb; zhang2017up; scholz_comparison_2021). However, the majority of these works do not focus specifically on IML models. While interpretable models are sometimes included in broader comparisons, recent state-of-the-art approaches such as SR (la2021contemporary), EBM (nori2019interpretml), or GOSDT (hu2019optimal; mctavish2022fast) are often omitted.
A recent IML-focused work, kruschel2025challenging, presents a systematic comparison of interpretable models with a focus on GAMs. Their analysis provides strong evidence against the presumed trade-off between accuracy and interpretability. Nevertheless, the scope of their study remains restricted, as other relevant interpretable approaches are not considered.
Moreover, most existing comparative evaluations, such as the work by fernandez2014we, rely on aggregated performance results across a collection of datasets, rather than systematically studying how model performance varies under different data conditions. As a result, important nuances in model behavior may be obscured. In contrast, scholz_comparison_2021 explicitly investigate how classification methods respond to varying structural dataset characteristics, such as dimensionality and target distributions, using synthetic data to evaluate performance under controlled conditions. Their investigation reveals that small heterogeneous ensembles, specifically simple averages of just 3 to 5 models, generally yield the best predictive performance. They further identify that specific settings favor certain algorithms: Bagged CART excels in low-dimensional, high-sample scenarios, while nearest shrunken neighbor classifiers are most effective for unbalanced datasets. This type of analysis underscores the importance of accounting for dataset properties in comparative studies, a perspective that remains largely unexplored in the context of IML. While synthetic datasets enable controlled experimentation, they may fail to fully capture the complexity, heterogeneity, and noise patterns found in real-world data, which can limit the external validity of the conclusions regarding which specific models are optimal for a given complexity scenario
@@
Another recurring limitation in many comparative evaluations is the restricted choice of benchmark datasets, often confined to a small number of manually selected examples, typically drawn from well-known repositories such as UCI (uci2019). This raises concerns regarding the generalizability and robustness of the reported findings. Moreover, rigorous statistical testing is frequently absent, further limiting the reliability of conclusions. As argued by herrmann2024position, empirical machine learning research is often presented as confirmatory, even though comparative evaluation should instead be treated as exploratory. Comparative analyses are also commonly conducted in the context of introducing new models, which can unintentionally bias results in favor of the proposed method.
In contrast to these practices, and to prior studies that either relied on narrow subsets of datasets (kruschel2025challenging) or on synthetic data designed to isolate specific conditions (scholz_comparison_2021), our evaluation will draw on a broad set of real-world datasets from PMLB, excluding only those unsuitable for our setting (see Section 3.2). This approach mitigates the risk of selection bias and ensures a more comprehensive assessment of model behavior across diverse problem types. Furthermore, to move beyond aggregate performance and gain insight into how models behave under different data conditions, we stratify results using several dataset-level metrics (see Section 5.1), enabling a more fine-grained and interpretable analysis and test generalization properties both in- and out-of-sample.
@@ 3 Experimental Setup
This section describes the experimental setup underlying our large-scale benchmark of IML models. We first present the set of candidate models included in the study and motivate their selection based on interpretability and implementation maturity. We then detail the dataset collection process and preprocessing pipeline applied to construct a consistent benchmark. Finally, we describe the computational environment and the hyperparameter search spaces and budgets used to ensure a fair and reproducible comparison across methods.
@@ 3.1 Selection of ML Models
Method Acronym Task Reference Library
Linear Regression LR REGR scikit-learn scikit-learn1
Logistic Regression LR CLF scikit-learn scikit-learn
Lasso Regression LASSO REGR scikit-learn scikit-learn
Generalized Linear Models GLM REGR scikit-learn scikit-learn
Polynomial Regression w/ LASSO PR+LASSO REGR scikit-learn scikit-learn
Explainable Boosting Machine EBM CLF, REGR nori2019interpretml InterpretML2
Interpretable Generalized Additive Neural Network IGANN CLF, REGR kraus2023interpretable IGANN3
Decision Trees DT CLF, REGR scikit-learn scikit-learn
Generalized Optimal Sparse Decision Trees GOSDT CLF mctavish2022fast GOSDT4
Naive Bayes NB CLF scikit-learn scikit-learn
k-Nearest Neighbors k-NN CLF, REGR scikit-learn scikit-learn
Symbolic Regression SR REGR cranmer2023interpretable PySR5
Table 1: Overview of the interpretable models included in our comparative evaluation. For each method, we report its acronym, the type of task it addresses (regression and/or classification), key references, and the software library or implementation used in our experiments.
@@
The baseline of our analysis is formed by the family of linear models (Linear/Logistic Regression, Lasso, and Polynomial Regression). While these methods offer the highest degree of transparency, their expressiveness is limited by their fixed functional form. Increasing the level of expresiveness, we also consider Generalized Additive Models. Within this class, we specifically selected Explainable Boosting Machines (nori2019interpretml) and Interpretable Generalized Additive Neural Networks (kraus2023interpretable). We focused on these two methods because recent comparative work by kruschel2025challenging positions them on the performance–interpretability Pareto frontier among modern GAM variants; furthermore, unlike many experimental alternatives, they offer mature and well-documented implementations.
To explore a further level of flexibility beyond the additive constraints of GAMs, we include Symbolic Regression. Unlike standard parametric or semi-parametric methods, SR does not assume a pre-defined model structure; instead, it searches the space of mathematical expressions to discover the underlying functional form of the data. Given that SR is computationally demanding, we limit our evaluation to a single representative: PySR (cranmer2023interpretable). We selected PySR over alternative implementations (such as Operon (operon) or QLattice (brolos2021approach)) as recent benchmarks demonstrate that it delivers highly competitive accuracy–interpretability trade-offs (de2024srbench++), alongside a mature and efficient Python implementation.
Finally, regarding tree-based methods, we include both classical greedy Decision Trees (DTs) and Generalized Optimal Sparse Decision Trees (mctavish2022fast). We include GOSDT because it employs a bound-driven search for provably optimal trees on binary variables. This approach results in smaller, more interpretable trees compared to greedy splitting methods used in standard DTs, while maintaining competitive predictive performance.
A detailed description and discussion of these methods is provided in Appendix A.

### v2

@@ Abstract
As machine learning models are increasingly deployed in high-stakes domains, the need for interpretability has grown to meet strict regulatory and accountability constraints. Despite this interest, systematic evaluations of inherently interpretable models for tabular data remain scarce and often focus solely on aggregated performance.
To address this gap, we evaluate sixteen interpretable methods, including Explainable Boosting Machines (EBMs), Symbolic Regression (SR), and Generalized Optimal Sparse Decision Trees, across 216 real-world tabular datasets. We assess predictive accuracy, computational efficiency, and generalization under distributional shifts. Moving beyond aggregate performance rankings, we further analyze how model behavior varies with dataset meta-features and operationalize these descriptors to study algorithm selection.
Our analyses reveal a clear dichotomy: in regression tasks, models exhibit a predictable performance hierarchy dominated by EBMs and SR that can be inferred from dataset characteristics. In contrast, classification performance remains highly dataset-dependent with no stable hierarchy, showing that standard complexity measures fail to provide actionable guidance. Furthermore, we identify an “interpretability tax”, showing that models explicitly optimizing for structural sparsity incur significantly longer training times. Overall, these findings provide practical guidance for practitioners seeking a balance between interpretability and predictive performance, and contribute to a deeper empirical understanding of interpretable modeling for tabular data.
@@ 1 Introduction
The field of Interpretable Machine Learning (IML) has emerged, emphasizing models whose internal mechanisms are transparent and understandable to human users [molnar2020interpretable]. This paradigm stands in contrast to the black box nature of complex architectures, such as deep neural networks, which prioritize predictive accuracy at the expense of explanatory insight [breiman2001statistical, rudin2019we].
@@
It is also necessary to distinguish intrinsic interpretability from post-hoc explainability (often termed XAI). Post-hoc methods attempt to explain a complex model after it has been trained; however, because these are merely approximations, they may not accurately reflect how the model actually made a decision. In contrast, intrinsically interpretable models are transparent by design, guaranteeing that the explanation is an exact representation of the model’s internal logic [rudin_stop_2019]. For this reason, such techniques are increasingly adopted in domains such as healthcare, finance, and law, where accountability and the understanding of the model decision-making process are essential [doshi2017towards, rudin_stop_2019]. This demand is further amplified by evolving regulatory frameworks, such as the European Union’s AI Act and the General Data Protection Regulation (GDPR), which mandate stringent transparency requirements and the right to explanation for high-risk automated systems. In many tabular-data applications, intrinsically interpretable models, such as linear models, sparse decision trees, or symbolic models, are already regarded as strong candidates due to their clarity and ease of validation by domain experts [zeng2017interpretable, la2021contemporary]. Despite this growing practical importance, empirical guidance on the comparative performance of interpretable models remains limited.
@@
Most existing benchmarking studies focus on general-purpose machine learning algorithms rather than interpretable ones [fernandez2014we, wainer2016comparison, olson2017pmlb, scholz_comparison_2021]. As a result, practitioners often lack systematic evidence about (i) whether some interpretable models tend to perform better than others across diverse real-world datasets, and (ii) how model performance changes as a function of dataset characteristics such as dimensionality, sample size, linearity, or class imbalance. A few recent works examine specific subclasses of interpretable models, such as Generalized Additive Models (GAMs) [kruschel2025challenging], but do not cover the broader landscape of modern IML approaches nor analyze performance variation under different data conditions. Moreover, comparative evaluations are typically conducted on small collections of hand-selected datasets or on synthetic data, which may limit generalizability.
@@
At the same time, the ecosystem of interpretable methods has expanded substantially over the past decade. Techniques such as Explainable Boosting Machines (EBMs) [nori2019interpretml], Interpretable Generalized Additive Neural Networks (IGANNs) [kraus2023interpretable], Generalized Optimal Sparse Decision Trees (GOSDTs) [mctavish2022fast], and modern Symbolic Regression methods [la2021contemporary] provide new modeling options whose empirical properties have not yet been systematically compared. Consequently, users are faced with an increasingly fragmented methodological landscape without clear, data-driven recommendations for model selection.
@@
  • Using the full Penn Machine Learning Benchmark (PMLB) repository [olson2017pmlb], we evaluate 16 interpretable models across 137 regression and 79 classification datasets. Our analysis includes both traditional and recent IML methods with stable, well-maintained implementations (9 regression methods and 7 classification methods), assessing both predictive performance and training time. Our experimental design includes 4-fold cross-validation, systematic hyperparameter optimization, and rigorous preprocessing, resulting in over 14,000 model evaluations and more than 700,000 individual training runs.
@@
  • Beyond aggregated results, we investigate how model performance varies with key dataset characteristics, also called meta-features, including structural properties (e.g., number of features and samples), degree of linearity, and class imbalance [maciel_measuring_2016, lorena_how_2020]. This stratified analysis provides deeper insights into the conditions under which different interpretable models excel or underperform, offering practical guidance for both researchers and practitioners.
  • Furthermore, we operationalize these meta-features to address the Algorithm Selection Problem. By training predictive meta-models, we quantify inherent dataset difficulty and estimate pairwise algorithm superiority, demonstrating the extent to which dataset characteristics can reliably guide a priori model selection across different domains.
While interpretability is a multifaceted concept, many of its aspects rely on subjective human evaluation [doshi2017towards, lipton2018mythos, arrieta2020explainable]. To enable a consistent large-scale empirical comparison, we focus on structural proxies of interpretability, such as sparsity in linear models and tree size in decision-based methods. Because interpretability is not directly comparable across heterogeneous model classes in a consistent and objective manner, we evaluate intrinsically interpretable methods primarily through their empirical effectiveness, i.e., their predictive performance across datasets with diverse characteristics, and analyze how this effectiveness varies with dataset meta-features.
Our empirical analysis uncovers a fundamental dichotomy between regression and classification tasks regarding both performance hierarchy and predictability. In the regression domain, we observe a clear and consistent stratification of models, where performance is strongly governed by the underlying geometric properties of the data. In contrast, classification exhibits no stable or recoverable hierarchy. Model performances are highly overlapping and strongly dataset-dependent, preventing the emergence of a consistent ranking across problems.
By operationalizing dataset meta-features, we further demonstrate that this contrast extends to algorithm selection. In regression, meta-features enable a reliable a priori prediction of model performance and pairwise superiority. In classification, however, these same descriptors fail to provide actionable guidance: neither dataset stratification nor meta-modeling yields reliable predictive power, highlighting a fundamental limitation of current data complexity measures in this setting. Furthermore, we quantify a pervasive computational trade-off inherent to interpretable modeling. We show that enforcing structural parsimony or capturing complex non-linear relationships often requires a significant increase in training time compared to standard baselines, a "tax" that practitioners must weigh against the benefits of transparency. Collectively, these findings provide a data-driven framework for balancing predictive accuracy, structural simplicity, and computational cost in high-stakes applications.
The remainder of this paper is organized as follows. Section 2 reviews related work on benchmarking, focusing on interpretable models. Next, Section 3 details the experimental setup, including model selection and dataset preprocessing. We then specify the evaluation protocol in Section 4, defining the cross-validation design, generalization scenarios, performance metrics, and meta-features used for stratification. Section 5 presents and analyzes the empirical results for both regression and classification tasks. Finally, Section 6 contextualizes these findings within the existing literature to offer practical guidelines, while Section 7 concludes the paper with directions for future research.
@@ 2 Related Work
A number of studies in the literature have conducted comparative analyses of machine learning models [fernandez2014we, wainer2016comparison, olson2017pmlb, zhang2017up, scholz_comparison_2021]. However, the majority of these works do not focus specifically on IML models. While interpretable models are sometimes included in broader comparisons, recent state-of-the-art approaches such as SR [la2021contemporary], EBM [nori2019interpretml], or GOSDT [hu2019optimal, mctavish2022fast] are often omitted.
A recent work focusing on IML, kruschel2025challenging, presents a systematic comparison of interpretable models with a focus on GAMs. Their analysis provides strong evidence against the presumed trade-off between accuracy and interpretability. Nevertheless, the scope of their study remains restricted, as other relevant interpretable approaches are not considered.
Moreover, most existing comparative evaluations, such as the work by fernandez2014we, rely on aggregated performance results across a collection of datasets, rather than systematically studying how model performance varies under different data conditions. As a result, important nuances in model behavior may be obscured. In contrast, scholz_comparison_2021 explicitly investigate how classification methods respond to varying structural dataset characteristics, such as dimensionality and target distributions, using synthetic data to evaluate performance under controlled conditions. Their investigation reveals that small heterogeneous ensembles generally yield the best predictive performance. They further identify that specific settings favor certain algorithms: Bagged CART excels in low-dimensional, high-sample scenarios, while nearest shrunken neighbor classifiers are most effective for unbalanced datasets. This type of analysis underscores the importance of accounting for dataset properties in comparative studies, a perspective that remains largely unexplored in the context of IML. While synthetic datasets enable controlled experimentation, they may fail to fully capture the complexity, heterogeneity, and noise patterns found in real-world data, which can limit the external validity of the conclusions regarding which specific models are optimal for a given complexity scenario.
@@
To systematically move beyond aggregate metrics and basic structural properties, our work draws inspiration from the meta-learning literature. Meta-learning addresses the Algorithm Selection Problem, building on the premise that no single learning algorithm is universally superior [khan2020literature]. Instead of seeking a universal winner, empirical meta-learning studies map expected algorithm performance to measurable characteristics of datasets, known as meta-features [vanschoren2018meta, rivolli2018characterizing, rivolli2022meta]. Existing results demonstrate that analyzing these meta-features allows researchers to accurately outline the specific “domains of competence" of different learning algorithms [lorena_how_2020].
Notably, data complexity measures, a specialized subset of meta-features originally introduced by ho2002complexity, have proven highly effective in this regard. By quantifying the geometrical and topological difficulty of problems (such as class overlap, linear separability, and neighborhood density), empirical studies have shown that these measures successfully predict when algorithms like nearest neighbors or specific linear classifiers will succeed or fail [lorena_how_2020]. Recently, these complexity measures have been successfully adapted to the regression domain, demonstrating their effectiveness in predicting regressor performance based on target correlation and function smoothness [lorena2018data].
Another recurring limitation in many comparative evaluations is the restricted choice of benchmark datasets, often confined to a small number of manually selected examples, typically drawn from well-known repositories such as UCI [uci2019]. This raises concerns regarding the generalizability and robustness of the reported findings. Moreover, rigorous statistical testing is frequently absent, further limiting the reliability of conclusions. As argued by herrmann2024position, empirical machine learning research is often presented as confirmatory, even though comparative evaluation should instead be treated as exploratory. Comparative analyses are also commonly conducted in the context of introducing new models, which can unintentionally bias results in favor of the proposed method.
In contrast to these practices, and to prior studies that either relied on narrow subsets of datasets [kruschel2025challenging] or on synthetic data designed to isolate specific conditions [scholz_comparison_2021], our evaluation will draw on a broad set of real-world datasets from PMLB, excluding only those unsuitable for our setting (see Section 3.2). This approach mitigates the risk of selection bias and ensures a more comprehensive assessment of model behavior across diverse problem types. Furthermore, to move beyond aggregate performance and gain insight into how models behave under different data conditions, we stratify results using several dataset-level metrics, or meta-features (see Section 4.4), enabling a more fine-grained and interpretable analysis and test generalization properties both in- and out-of-sample.
@@ 3 Experimental Setup
This section describes the experimental setup underlying our large-scale benchmark of IML models. We first present the set of candidate models included in the study and motivate their selection. We then detail the dataset collection process and preprocessing pipeline applied to construct a consistent benchmark. Finally, we describe the computational environment and the hyperparameter search spaces and budgets used to ensure a fair and reproducible comparison across methods.
@@ 3.1 Selection of ML Models
Method Acronym Task Type Reference Library
Linear Regression LR REGR DEC scikit-learn scikit-learn1
Logistic Regression LR CLF DEC scikit-learn scikit-learn
Lasso Regression LASSO REGR DEC scikit-learn scikit-learn
Generalized Linear Models GLM REGR DEC scikit-learn scikit-learn
Polynomial Regression w/ LASSO PR+LASSO REGR DEC scikit-learn scikit-learn
Explainable Boosting Machine EBM CLF, REGR DEC nori2019interpretml InterpretML2
Interpretable Generalized Additive Neural Network IGANN CLF, REGR DEC kraus2023interpretable IGANN3
Decision Trees DT CLF, REGR SIM scikit-learn scikit-learn
Generalized Optimal Sparse Decision Trees GOSDT CLF SIM mctavish2022fast GOSDT4
Naive Bayes NB CLF ALG scikit-learn scikit-learn
k-Nearest Neighbors k-NN CLF, REGR ALG scikit-learn scikit-learn
Symbolic Regression SR REGR SIM cranmer2023interpretable PySR5
Table 1: Overview of the interpretable models included in our comparative evaluation. For each method, we report its acronym, the type of task it addresses (regression and/or classification), its primary interpretability dimension (Simulatability [SIM], Decomposability [DEC], or Algorithmic Transparency [ALG]), key references, and the software library or implementation used in our experiments.
@@
The baseline of our analysis is formed by the family of linear models (Linear/Logistic Regression, Lasso, and Polynomial Regression). While these methods offer the highest degree of transparency, their expressiveness is limited by their fixed functional form. Increasing the level of expressiveness, we also consider Generalized Additive Models. Within this class, we specifically selected Explainable Boosting Machines [nori2019interpretml] and Interpretable Generalized Additive Neural Networks [kraus2023interpretable]. We focused on these two methods because recent comparative work by kruschel2025challenging positions them on the performance–interpretability Pareto frontier among modern GAM variants; furthermore, unlike many experimental alternatives, they offer mature and well-documented implementations.
To explore a further level of flexibility beyond the additive constraints of GAMs, we include Symbolic Regression [kronberger2024symbolic]. Unlike standard parametric or semi-parametric methods, SR does not assume a pre-defined model structure; instead, it searches the space of mathematical expressions to discover the underlying functional form of the data. Given that SR is computationally demanding, we limit our evaluation to a single representative: PySR [cranmer2023interpretable]. We selected PySR over alternative implementations (such as Operon [operon] or QLattice [brolos2021approach]) as recent benchmarks demonstrate that it delivers highly competitive accuracy–interpretability trade-offs [de2024srbench++], alongside a mature and efficient Python implementation.
Finally, regarding tree-based methods, we include both classical greedy Decision Trees (DTs) and Generalized Optimal Sparse Decision Trees [mctavish2022fast]. We include GOSDT because it employs a bound-driven search for provably optimal trees on binary variables. This approach results in smaller, more interpretable trees compared to greedy splitting methods used in standard DTs, while maintaining competitive predictive performance.
Altogether, our selection of these methods ensures a comprehensive evaluation that spans three core dimensions of interpretability: simulatability, decomposability, and algorithmic transparency. For a detailed discussion of each individual method, as well as a taxonomy mapping these models to their respective interpretability dimensions, we refer the reader to Appendices A and B.